# Small but large: Single organic molecules as hybrid platforms for quantum technologies

### v1

@@ II.2.1 Intermolecular Vibrational States
The typical vibrational landscape of a molecular host-guest system is shown with an example of a perylene-doped naphthalene and anthracene crystal in Fig. 3(a), where the phonon density of states in naphthalene crystal (red lines), anthracene crystal (black lines) and intramolecular vibrations of a perylene molecule (green lines) are plotted up to $12$ THz. Among many mechanical degrees of freedom, phonons corresponding to the center-of-mass translational motion of host molecules are the lowest in energy scale. These modes approximately span the frequency range $0-4.5$ THz with a cut-off frequency determined by the unit-cell size of the host molecular crystal, i.e., around $1$ nm for PAH molecular crystals. Assuming molecules as rigid bodies, there are $63N-$ 6 lattice modes for $N$ number of molecules in the crystal, which practically renders the spectrum of such modes as a continuum. These modes can be classified as acoustic and optical phonons based on their properties at the zero-phonon wavevector. Acoustic phonons result from in-phase motion of molecules as rigid bodies and their frequencies range in between $0-1.5$ THz for PAH molecular crystals [see Fig. 3(b)]. These modes possess linear dispersion for low frequencies, i.e., up to $\sim 100-200$ GHz, which results in a small density of states. The lifetime of these modes is determined by absorption and boundary scattering effects and can reach up to millisecond timescales for modes up to megahertz frequencies [52]. However, the electron-phonon couplings to these modes are small due to large phonon mode volumes in the molecular crystal [38]. Hence, optically accessing these modes is difficult.
@@ II.2.2 Intramolecular Vibrational States
The molecular host-guest system additionally possesses intramolecular vibrations of the guest molecule [see Fig. 3(a)]. There are $63n-$ 6 vibrations corresponding to $n$ number atoms in the molecule, and they correspond to librational and bond vibrations of the guest molecule, reaching up to $93$ THz for the C-H stretching motion. Frequencies of librational modes are determined by the size of the molecule. The lowest libration frequency for perylene molecule in vacuum is $0.9$ THz. In the range of $0-5$ THz, librational modes of the guest molecule hybridize with the translational lattice modes. This results in slightly delocalized vibrational states over the host molecules, so-called pseudolocal phonon modes [67] [see Fig. 3(c)]. Depending on the insertion of the guest molecule, some of these modes are Raman (Franck-Condon) active, and due to the semi-localized nature of their vibration over the guest molecule, they can have large electron-vibron coupling strengths compared to phonon modes of the host matrix. This is evident in the single molecule emission spectra, where the phonon-wing spectrum shows pronounced Lorentzian peaks below $5$ THz with lifetimes in the order of picoseconds [68]. In addition, these modes are shown to be responsible for temperature-dependent line broadening of the ZPL transition [54].
The rest of the intramolecular vibrations are localized on the guest molecule as the host vibrational density of states is sparse [see Figs. 3(a, c)]. Hence, these modes have very large electron-vibron coupling strengths in the order of their frequencies [46, 69, 42]. Such ultra-strong couplings can be witnessed from the emission/absorption spectrum of the host-guest system.
@@ II.2.3 Vibrational Relaxation
Despite long radiative lifetimes of molecular vibrations in the gas phase that range up to seconds [70, 71], the nonradiative vibrational relaxation in a solid typically yields lifetimes in the order of picosecond irrespective of the frequency and temperature range [69]. This process stems from the anharmonicity of the potential energy surface along host and guest vibrational coordinates, and results in short vibrational lifetimes even though certain vibrations of the guest molecule are in principle mechanically shielded due to gaps in the host vibrational band [see Fig. 3(a)]. Based on the perturbative treatment of this problem, the vibrational relaxation process due to cubic anharmonic interactions between vibrational modes can be divided into three categories depending on the frequency of the guest vibrational mode ($\omega_{\nu_{v}}$[72], as shown in Fig. 3(d). In the first process, the guest molecule’s vibrational frequency is smaller than two times the phonon cut-off frequency ($\omega_{\text{max}}$), i.e., cut-off frequency of the two-phonon density of states, such that the vibration of the guest molecule relaxes via two lattice phonons based on energy and momentum conservation as a result of the three-body interaction between the normal-mode displacements. The second process happens when the guest molecule’s vibrational frequency lies above the cut-off frequency of the two-phonon density of states, where an intramolecular vibration of the guest molecule could assist the non-radiative relaxation. If the guest molecule’s vibrational frequency is too large compared to the phonon cut-off frequency, the vibrational relaxation occurs via intramolecular relaxation in the guest molecule as the third process. Besides, as this picture is perturbative, the higher-order anharmonicities between vibrational modes also play a role once the energy conservation cannot be satisfied via relaxation to two vibrations. As a result, despite the weak-couplings between the intramolecular guest vibrations and phonon modes, the immense number of decay paths provided by the continuum of phonon modes yield picosecond-scale vibrational lifetimes. Despite such a short lifetimes, these vibrations are mostly in their quantum ground state even at room temperature, an interesting property that could be exploited to push quantum technologies to room temperature [34, 73].
@@ II.3 Spin states
The spin states of electrons and nuclei in the molecular host-guest system offer additional parameter spaces that hold potential for quantum applications. We start by introducing the properties of the lowest triplet state |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ and its role in the photophysics of the molecule.
@@
In the |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ state, the two unpaired electrons interact through the magnetic dipole-dipole interaction, which can be described by the phenomenological model Hamiltonian $\mathbf{H}_{\text{ZFS}}=\mathbf{S}\cdot\mathbf{D}\cdot\mathbf{S}$, with $\mathbf{S}$ the total spin operator and $\mathbf{D}$ the fine-structure tensor [74]. For PAH molecules with $D_{2h}$ symmetry, such as pentacene and perylene, $\mathbf{D}$ can be diagonalized in the coordinate system defined by the geometrical symmetry axes $\alpha=x,y,z$ of the molecule, as illustrated with a perylene molecule in Fig. 4(a). The three eigenstates associated with the three principal vectors, known as the zero-field eigenstates, can be labeled |$\ket{T_{1}^{\alpha}}$. In |$\ket{T_{1}^{\alpha}}$, the magnetic moment of the electron spin is oriented in the plane perpendicular to the axis $\alpha$. As shown in Fig. 4(b), the ZFS is characterized by two parameters $D$ and $E$, which correlate with the delocalization of the spin density in the $\pi$ orbital and the axial symmetry of the molecule, respectively. Transitions between any pair of the zero-field states are allowed by magnetic-dipole interactions and can be driven by microwave fields [74].
@@
In the triplet manifolds, the electron spin couples to nuclear spin wavefunction $\phi_{\text{nuc}}(\mathbf{I})$ through the hyperfine interaction. This interaction is described by the Hamiltonian $\mathbf{H}_{\text{HFI}}=\sum_{i}\mathbf{S}\cdot{\mathbf{A}}^{(i)}\cdot{\mathbf{I}_{i}}$, where $\mathbf{I}_{i}$ is the nuclear magnetic moment of the $i^{\text{th}}$ nucleus, and $\mathbf{A}^{(i)}$ is its hyperfine interaction tensor with the electron spin. In hydrocarbon molecular crystals, nuclear spins mainly originate from protons and 13C isotopes in the guest molecule and the host matrix. Notably, protons can be replaced by deuterons via isotope exchange reactions, through which the number of spin-1/2 nuclei in the guest molecule and the matrix can be significantly reduced. Using deuterated matrices makes it possible to resolve the hyperfine splitting due to individual nuclear spins in the molecule [75, 76]. The right panel of Fig. 4(b) illustrates the hyperfine splitting with a magnetic field applied along the $z$ axis and considering only a single spin-$1/2$ nucleus in a perylene molecule. Similar to the demonstration with nitrogen-vacancy centers in diamond [77], hyperfine splitting offers the opportunity to selectively access individual nuclear spin states using narrow-band microwave pulses, and facilitate the realization of controlled-rotation gates between electron and nuclear spins [77].
@@
Figure 4: (a) Structure of a perylene molecule and its three symmetry axes. (b) Left: Zero-field splitting of the three triplet sublevels |$\ket{T_{1}^{\alpha}}$ in the absence of an external magnetic field ($\mathbf{B}=0$), characterized by parameters $D$ and $E$. Right: Zeemann shift and hyperfine splitting due to a spin-$1/2$ nuclear spin, when an external magnetic field $B_{z}$ is applied along the $z$ axis. In the regime of a weak magnetic field, the zero-field bases |$\ket{T_{1}^{\alpha}}$ are still approximated eigenstates of the electron spin and $\ket{\uparrow}$, $\ket{\downarrow}$ denote the two states of the nuclear spin. Microwave (MW) driving of the electron spin enables controlled-rotation gates, conditioned on the state of the nuclear spin.
@@
The triplet state can be populated through optical pumping on the ZPL transition. From the |$\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ state, the molecule can occasionally cross over to a vibrationally excited level in the triplet manifold |$\ket{T_{j};\bm{\nu}_{v}\neq\bm{0},\bm{\nu}_{p}\neq\bm{0}}$. This process, known as intersystem crossing (ISC), is weakly allowed due to spin-orbital coupling, and typically assisted by intermolecular and intramolecular vibrations [78]. Following ISC, the molecule undergoes nonradiative relaxation and reaches the lowest triplet state |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$. The rate of ISC also depends on the final spin sublevel |$\ket{T_{1}^{\alpha}}$ of the triplet manifold [44]. From the state |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$, the molecule can decay to the ground state manifold either via phosphorescence, or through ISC to a |$\ket{S_{0};\bm{\nu}_{v}\neq\bm{0},\bm{\nu}_{p}\neq\bm{0}}$ level followed by vibrational relaxation. These two processes compete with each other. The rate of ISC has been shown to follow the energy-gap law [79], while its underlying microscopic mechanisms are generally complex [80]. Since both processes are spin-forbidden, the lifetime of the |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ state is long, reaching up to seconds [81].
@@
In the triplet state, the molecule does not fluoresce, so a decrease in fluorescence intensity can serve as an indicator of triplet-state occupation. By using a resonant microwave drive to shuffle the population between two of the three sub-levels, the mean residence time in the triplet state can be altered, resulting in changes to the steady-state fluorescence intensity. This technique was applied to pentacene molecules in $para$-terphenyl crystal in the 1990s, which led to the first detections of electron spin transitions in a single molecule [82, 83]. These efforts were later extended to deuterated pentacene molecules and enabled the spectroscopic detection of the hyperfine splittings due to individual 13C and proton spins [75, 76]. Very recently, the detection and coherent control of triplet electron spins have been demonstrated on an ensemble of pentacene molecules at room temperature [39], potentially setting the ground for ambient quantum sensing using molecular host-guest materials.
@@ III Perspectives in Quantum Technology
So far, single organic molecules have found applications as single photon sources [6, 7], quantum optical nonlinear elements [8, 9, 10], and nanoscale sensors for probing local field properties in materials [15, 31]. They have also been integrated into nanophotonic and hybrid nanostructures, such as nanoantennas, waveguides and optical resonators, laying out promising future paths in photonic quantum technologies [3]. The vibrational and spin states of molecular host-guest systems, as discussed in the previous section, hold significant potential for expanding the application of these systems. Furthermore, the chemical diversity of molecules could enhance this potential by providing molecular candidates with tailored quantum properties. In the following, we outline possible strategies to harness the vibrational and spin states for applications that have not been realized on these systems before and discuss the opportunities presented by the vast chemical space.
@@ III.1 Vibrational quantum memory
The schematic of a vibrational quantum memory in a single molecule is shown in Fig. 5, where the molecular electronic transition provides the interface to flying qubits and quantum memory is realized in the ground state vibrational manifold. The flying qubits encoded in photon-number [85] or time-bin degrees of freedom [86] can be stored and retrieved with well-established techniques such as stimulated Raman scattering [87, 38] as shown in Fig. 5 or stimulated Raman adiabatic passage [88] or heralded Stokes Raman scattering [85]. Despite the sizable couplings between the electronic and vibrational levels, the picosecond-timescale vibrational lifetimes in these systems limit such applications. Although the vibrational relaxation mechanism is challenging to control in the solid state, it has recently been proposed that low-frequency acoustic phonon modes in the host crystal can be harnessed as quantum memory elements once the molecular environment is engineered [38]. These so-called continuum modes can be made spectrally resolved by employing nanometer-sized host crystals, compatible with state-of-the-art fabrication techniques [89, 90]. In addition, the lifetimes of these phonon modes can be made long-lived up to milliseconds via phononic crystal engineering techniques. Although such a quantum memory operates at millikelvin temperatures due to the low phonon frequencies involved, the higher frequency vibrations such as optical phonons in the host matrix and intramolecular vibrations of the guest molecule could be utilized to realize quantum memory elements operating at high temperatures once their decay paths are controlled. The latter could be achieved by choosing host materials with a sparse vibrational density of states such as benzene or by placing small molecules on a surface vibrationally decoupled from the substrate beneath [91].

### v2

@@ II.2.1 Intermolecular Vibrational States
The typical vibrational landscape of a molecular host-guest system is shown with an example of a perylene-doped naphthalene and anthracene crystal in Fig. 3(a), where the phonon density of states in naphthalene crystal (red lines), anthracene crystal (black lines) and intramolecular vibrations of a perylene molecule (green vertical lines) are plotted up to $12$ THz. Among many mechanical degrees of freedom, phonons corresponding to the center-of-mass translational motion of host molecules are the lowest in energy scale. These modes approximately span the frequency range $10$ kHz - $4.5$ THz with a cut-off frequency determined by the unit-cell size of the host molecular crystal, i.e., around $1$ nm for PAH molecular crystals. Assuming molecules as rigid bodies, there are $63N-$ 6 lattice modes for $N$ number of molecules in the crystal, which practically renders the spectrum of such modes as a continuum. These modes can be classified as acoustic and optical phonons based on their properties at the zero-phonon wavevector. Acoustic phonons result from in-phase motion of molecules as rigid bodies [see Fig. 3(b)] and their frequencies range in between $0-1.5$ THz for PAH molecular crystals. These modes possess linear dispersion for low frequencies, i.e., up to $\sim 100-200$ GHz, which results in a small density of states. The lifetime of these modes is determined by absorption and boundary scattering effects and can reach up to millisecond timescales for modes up to megahertz frequencies [54]. However, the electron-phonon couplings to these modes are small due to large phonon mode volumes in the molecular crystal [40]. Hence, optically accessing these modes is difficult.
@@ II.2.2 Intramolecular Vibrational States
The molecular host-guest system additionally possesses intramolecular vibrations of the guest molecule [see the lower panel of Fig. 3(a)]. There are $63n-$ 6 vibrations corresponding to $n$ number atoms in the molecule, and they correspond to librational and bond vibrations of the guest molecule, reaching up to $93$ THz for the C-H stretching motion. Frequencies of librational modes are determined by the size of the molecule. The lowest libration frequency for perylene molecule in vacuum is $0.9$ THz. In the range of $0-5$ THz, librational modes of the guest molecule hybridize with the translational lattice modes. This results in slightly delocalized vibrational states over the host molecules [see Fig. 3(c)], so-called pseudolocal phonon modes [71]. Depending on the insertion of the guest molecule, some of these modes are Raman (Franck-Condon) active, and due to the semi-localized nature of their vibration over the guest molecule, they can have large electron-vibron coupling strengths compared to phonon modes of the host matrix. This is evident in the single molecule emission spectra, where the phonon-wing spectrum shows pronounced Lorentzian peaks below $5$ THz with lifetimes in the order of picoseconds [72]. In addition, these modes are shown to be responsible for temperature-dependent line broadening of the ZPL transition [58].
The rest of the intramolecular vibrations are localized on the guest molecule [see Figs. 3(c)] as the host vibrational density of states is sparse. Hence, these modes have very large electron-vibron coupling strengths in the order of their frequencies [57, 44]. Such ultra-strong couplings can be witnessed from the emission and absorption spectrum of the host-guest system.
@@ II.2.3 Vibrational Relaxation
Despite long radiative lifetimes of molecular vibrations in the gas phase that range up to seconds [73, 74], the nonradiative vibrational relaxation in a solid typically yields lifetimes in the order of picosecond irrespective of the frequency and temperature range [57]. This process stems from the anharmonicity of the potential energy surface along host and guest vibrational coordinates, and results in short vibrational lifetimes even though certain vibrations of the guest molecule are in principle mechanically shielded due to gaps in the host vibrational band, as seen within the $7-11$ THz spectral region of the phonon density of states and perylene vibrational spectrum in Fig. 3(a). Based on the perturbative treatment of this problem, the vibrational relaxation process due to cubic anharmonic interactions between vibrational modes can be divided into three categories [75] depending on the frequency of the guest vibrational mode, $\omega_{\nu_{v}}$, as shown in Fig. 3(d). In the first process, the guest molecule’s vibrational frequency is smaller than two times the phonon cut-off frequency, $\omega_{\text{max}}$, i.e., cut-off frequency of the two-phonon density of states, such that the vibration of the guest molecule relaxes via two lattice phonons based on energy and momentum conservation as a result of the three-body interaction between the normal-mode displacements. The second process happens when the guest molecule’s vibrational frequency lies above the cut-off frequency of the two-phonon density of states, where an intramolecular vibration of the guest molecule could assist the non-radiative relaxation. If the guest molecule’s vibrational frequency is too large compared to the phonon cut-off frequency, the vibrational relaxation occurs via intramolecular relaxation in the guest molecule as the third process. Besides, as this picture is perturbative, the higher-order anharmonicities between vibrational modes also play a role once the energy conservation cannot be satisfied via relaxation to two vibrations. As a result, despite the weak couplings between the intramolecular guest vibrations and phonon modes, the immense number of decay paths provided by the continuum of phonon modes yield picosecond-scale vibrational lifetimes. Even with such short lifetimes, these vibrations are mostly in their quantum ground state even at room temperature, an interesting property that could be exploited to push optomechanical quantum technologies beyond millikelvin temperatures [37].
@@ II.3 Spin states
The spin states of electrons and nuclei in the molecular host-guest system offer additional parameter spaces that hold potential for quantum applications. We start by introducing the properties of the lowest triplet state |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ and its role in the photophysics of molecules.
@@
In the |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ state, the two unpaired electrons interact through magnetic dipole-dipole interactions, which can be described by the phenomenological model Hamiltonian $\mathbf{H}_{\text{ZFS}}=\mathbf{S}\cdot\mathbf{D}\cdot\mathbf{S}$, with $\mathbf{S}$ the total spin operator and $\mathbf{D}$ the fine-structure tensor [76]. For PAH molecules with $D_{2h}$ symmetry, such as pentacene and perylene, $\mathbf{D}$ can be diagonalized in the coordinate system defined by the geometrical symmetry axes $\alpha=x,y,z$ of the molecule, as illustrated with a perylene molecule in Fig. 4(a). The three eigenstates associated with the three principal vectors, known as the zero-field eigenstates, can be labelled |$\ket{T_{1}^{\alpha}}$. In |$\ket{T_{1}^{\alpha}}$, the magnetic moment of the electron spin is oriented in the plane perpendicular to the axis $\alpha$. As shown in Fig. 4(b), the zero-field splitting is characterized by two parameters $D$ and $E$, which correlate with the delocalization of the spin density and the axial symmetry of the molecule, respectively. Transitions between any pair of the zero-field states are allowed by magnetic-dipole interactions and can be driven by microwave fields [76].
@@
In the triplet manifolds, the electron spin couples to nuclear spin wavefunction $\phi_{\text{nuc}}(\mathbf{I})$ through the hyperfine interaction. This interaction is described by the Hamiltonian $\mathbf{H}_{\text{HFI}}=\sum_{i}\mathbf{S}\cdot{\mathbf{A}}^{(i)}\cdot{\mathbf{I}_{i}}$, where $\mathbf{I}_{i}$ is the nuclear magnetic moment of the $i^{\text{th}}$ nucleus, and $\mathbf{A}^{(i)}$ is its hyperfine interaction tensor with the electron spin. In hydrocarbon molecular crystals, nuclear spins mainly originate from protons and 13C isotopes in the guest molecule and the host matrix. Notably, protons can be replaced by deuterons via isotope exchange reactions, through which the number of spin-1/2 nuclei in the guest and the host can be significantly reduced. Using deuterated matrices makes it possible to resolve the hyperfine splitting due to individual nuclear spins in a single molecule [77, 78]. The right panel of Fig. 4(b) illustrates the hyperfine splitting with a magnetic field applied along the $z$ axis and considering only a single spin-$1/2$ nucleus in a perylene molecule. Similar to the demonstration with nitrogen-vacancy centers in diamond [79], hyperfine splitting offers the opportunity to selectively access individual nuclear spin states using narrow-band microwave pulses, and facilitate the realization of controlled-rotation gates between electron and nuclear spins. Note that in the case of nitrogen-vacancy centers, having a triplet ground state allows easy access to nuclear spin states. In PAH molecules, hyperfine interaction is only present in the triplet excited states. This feature increases the complexity of accessing the nuclear spins, but at the same time offers an optically controllable switch of the hyperfine interaction.
@@
Figure 4: (a) Structure of a perylene molecule and its three symmetry axes. (b) Left: Zero-field splitting of the three triplet sublevels |$\ket{T_{1}^{\alpha}}$ in the absence of an external magnetic field ($\mathbf{B}=0$), characterized by parameters $D$ and $E$. Right: Zeeman shift and hyperfine splitting due to a spin-$1/2$ nuclear spin, when an external magnetic field $B_{z}$ is applied along the $z$ axis. In the regime of a weak magnetic field, the zero-field bases |$\ket{T_{1}^{\alpha}}$ are still approximated eigenstates of the electron spin and $\ket{\uparrow}$, $\ket{\downarrow}$ denote the two states of the nuclear spin. Microwave (MW) driving of the electron spin enables controlled-rotation gates, conditioned on the state of the nuclear spin.
@@
The triplet state can be populated through optical pumping on the ZPL transition. From the |$\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ state, the molecule can occasionally cross over to a vibrationally excited level in the triplet manifold |$\ket{T_{j};\bm{\nu}_{v}\neq\bm{0},\bm{\nu}_{p}\neq\bm{0}}$. This process, known as intersystem crossing (ISC), is weakly allowed due to spin-orbital coupling, and typically assisted by intermolecular and intramolecular vibrations [80]. Following ISC, the molecule undergoes nonradiative relaxation and reaches the lowest triplet state |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$. The rate of ISC also depends on the final spin sublevel |$\ket{T_{1}^{\alpha}}$ of the triplet manifold [46]. From the state |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$, the molecule can decay to the ground state manifold either via phosphorescence, or through ISC to a |$\ket{S_{0};\bm{\nu}_{v}\neq\bm{0},\bm{\nu}_{p}\neq\bm{0}}$ level followed by vibrational relaxation. These two processes compete with each other. The rate of ISC has been shown to follow the energy-gap law [81], while its underlying microscopic mechanisms are generally complex [82]. Since both processes are spin-forbidden, the lifetime of the |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ state is long, reaching microseconds to milliseconds [83, 84].
@@
In the triplet state, the molecule does not fluoresce, so a decrease in fluorescence intensity can serve as an indicator of triplet-state occupation. By using a resonant microwave drive to shuffle the population between two of the three sublevels, the mean residence time in the triplet state can be altered, resulting in changes to the steady-state fluorescence intensity. This technique was applied to pentacene molecules in $para$-terphenyl crystal in the 1990s, which led to the first detections of electron spin transitions in a single molecule [85, 86]. These efforts were later extended to deuterated pentacene molecules and enabled the detection of the hyperfine splittings due to individual 13C and proton spins [77, 78]. Apart from fluorescence detection, spin transitions among the triplet sublevels have also been measured through atomic force microscopy on a single pentacene molecule [87]. Very recently, the detection and coherent control of triplet electron spins have been demonstrated on an ensemble of pentacene molecules at room temperature [41], potentially setting the ground for ambient quantum sensing using molecular host-guest materials.
@@ III Perspectives in Quantum Technology
So far, single organic molecules have found applications as single photon sources, quantum optical nonlinear elements, and nanoscale sensors for probing local field properties in materials. They have also been integrated into nanophotonic and hybrid nanostructures, such as nanoantennas, waveguides and optical resonators, laying out promising future paths in photonic quantum technologies [3]. The vibrational and spin states of molecular host-guest systems, as discussed in the previous section, hold significant potential for expanding the application of these systems. Furthermore, the chemical diversity of molecules could enhance this potential by providing molecular candidates with tailored quantum properties. In the following, we outline possible strategies to harness the vibrational and spin states for applications that have not been realized on these systems before and discuss the opportunities presented by the vast chemical space.
@@ III.1 Vibrational quantum memory
The schematic of a vibrational quantum memory in a single molecule is shown in Fig. 5, where the molecular electronic transition provides the interface to flying qubits and quantum memory is realized in the ground state vibrational manifold. The flying qubits encoded in photon-number [89] or time-bin degrees of freedom [90] can be stored and retrieved with well-established techniques such as stimulated Raman scattering [91, 40] as shown in Fig. 5, stimulated Raman adiabatic passage [92] or heralded Stokes Raman scattering [89].
Although the electronic and vibrational levels of molecules exhibit sizable couplings, the short vibrational lifetimes, which are typically on the picosecond timescale, pose a significant limitation for quantum memory applications. To address this challenge, it has been proposed that GHz-frequency acoustic phonon modes in the host crystal can be exploited as long-lived quantum memory elements [40], provided that the molecular environment is carefully engineered. These so-called continuum modes can be made spectrally resolvable by employing nanometer-scale host crystals, which are compatible with current nanofabrication techniques [93, 94]. Such small mode volumes further allow for strong electron-phonon coupling strengths that can exceed the decoherence rates of the excited electronic state. Moreover, lifetimes of these acoustic phonon modes can be extended up to the millisecond regime through phononic crystal engineering [95], enabling their use as vibrational quantum memories.
The operation of such a memory is inherently limited to millikelvin temperatures due to thermal phonon occupation but could potentially be extended to higher temperatures by employing bulk acoustic wave resonators [96]. Efficient operation additionally requires near-perfect coupling of single photons to the single molecule. While free-space optics can achieve up to $13\%$ coupling efficiency [12], photonic structures offer values approaching $93\%$ [22]. Since the quantum memory protocol relies on coherent interactions, both the quantum efficiency of the single molecule and dephasing of the phonon mode limit the achievable fidelity. While single-molecule emitters have typically high quantum efficiency, the fidelity of quantum memory may nonetheless be constrained by the dephasing of phonon modes, which remains insufficiently characterized.
In addition to the acoustic phonon modes, quasi-localized lattice vibrational modes [97] and two-level system defects [98] in host crystals are promising platforms for quantum memory implementation [99, 100]. These excitations exhibit strong coupling to electronic transitions and possess long relaxation times, often reaching up to the millisecond regime, which can be further enhanced at millikelvin temperatures or through phonon engineering techniques [101].
Considering the availability of nanophotonic interfaces, phonon engineering strategies, and the high quantum efficiencies achievable with single molecules, this platform presents a compelling route towards realizing molecular quantum memories.